OpenStudy (anonymous):

Surface Area: All edges of a cube are expanding at a rate of 6 centimeters per second. How fast is the surface area changing when each edge is (A) 2 centimeters and (B) 10 centimeters?

OpenStudy (anonymous):

Given: de/dt where e= length of an edge and t= time, find da/dt where a= surface area.

OpenStudy (anonymous):

We know that da/dt = de/dt * da/de

OpenStudy (anonymous):

The equation of surface area for a cube is \[e^{2}*6=a\]

OpenStudy (anonymous):

Once finding da/de for the above equation, plug it into the first equation with with the given values and you should be able to solve for da/dt :)

OpenStudy (anonymous):

Thank you. I utilized the given information along with your explanation, and received the following answers by working out the word problem. I hope the answers are correct. \[Part A: \frac{ da }{ dt }(2) = (12)(6)(2) = 144cm ^{3}/\sec\] \[Part B: \frac{ da }{ dt }(10) = (12)(6)(10) = 720cm ^{3}/\sec\]

OpenStudy (anonymous):

One second, I'll check them.

OpenStudy (anonymous):

Yes that matches with what I got. :) Good job.

OpenStudy (anonymous):

Thank you!! :)